OpenStudy (anonymous):

While simplifying some math work, Peter wrote on his paper that x3 • x3 • x3 • x3 equaled x3+ 3 + 3 +3 . Did Peter simplify his work correctly and completely to a final answer? Would Peter’s work be the same if he were to simplify x3 + x3 + x3 + x3?

OpenStudy (anonymous):

@E.ali

OpenStudy (anonymous):

no peters simplification is wrong

OpenStudy (anonymous):

It is?

OpenStudy (anonymous):

a^b*a^c=a^(b+c) this is not equal to a^b+a^c

OpenStudy (anonymous):

peters simplification is right in the first case but the latter one is wrong i meant

OpenStudy (anonymous):

but it looks like the simplification in question is\[\large x^3\cdot x^3\cdot x^3\cdot x^3=x^{3+3+3+3}\]

OpenStudy (anonymous):

ammm... But I think is ... because computers have another langouege :) do u know any thinks about radix ?

OpenStudy (anonymous):

the above simplification is right

OpenStudy (anonymous):

radix is nothing different from base of a number system it informs about the structural design of a number system and conveys the number of digits used in that system

OpenStudy (yttrium):

Yes is the answer for the first question since it will arrive at x^12. But in the Second (addition) it will absolutely not applicable. Because that adding of exponents is only applicable in multiplication of same base. (You see they are all the same x that is why you can add all of the exponent) But in addition, if you have same term, what you can only do is to add the coefficient. Since they are all x^3 you can add them up. Arriving at 4x3. Do you understand me? Sorry for very long answer

OpenStudy (anonymous):

dont be sorry for a long answer if it helps :)

OpenStudy (anonymous):

but how does x3 • x3 • x3 • x3 equaled x3+ 3 + 3 +3 and x3 + x3 + x3 + x3= 4x3

OpenStudy (anonymous):

x3 + x3 + x3 + x3= 4x3 makes sense but im confused about the x3 • x3 • x3 • x3 equaling x3+ 3 + 3 +3

OpenStudy (anonymous):

They are all exponents btw

OpenStudy (yttrium):

It will arrive at \[x ^{3+3+3+3} = x ^{12}]

OpenStudy (anonymous):

Thank you

OpenStudy (yttrium):

No problem @bbbbbbbbbbbbbbbbbb

OpenStudy (anonymous):

The information included in this question is the intellectual property of Florida Virtual School. Posting questions from any FLVS course and receiving answers is considered cheating. When you are having help with your work, all you need to do is call your instructor - they are there to help you every step of the way!